Find the Laplace Transform of: (1+2at)t^(-1/2)exp(at)
Answer #1

fit) = (1+zat) + V2 gat flt) = (+42 + 2at +-12) pat f(t) = 72 pat +2a+ V pat 2482t)) = 244-72 pat + 2a + pat) E két-1/2 pat

f(s) 2(f(t)) = ef from * . (- ی) te ) ) = F ها تلد = ( هم )1( 2 اره-ی) from ** - ق ال : (4 ممال و فرم ها

برج اليه + T UH ) = )1 JT و + 42 (ه-ی) 2/ ( S) (4+ 4-ی لاری ST 3 / 2 ری-ک) لا رود = ( ه م - (غوا ]11